OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through the point (4, -1) and is perpendicular to the line whose equation is 2x - y - 7 = 0?

OpenStudy (anonymous):

@ChemGrey @amistre64

OpenStudy (anonymous):

line given is 2x - y - 7 = 0. in y = mx + c form => y = 2x - 7 so m1 = 2 for a line perpendicular to this, m = -1/m1 so m = -1/2 in slope intercept form.. (y - y1) = m(x - x1) x1 = 4, y1 = -1 so equation is => (y +1) = -1(x - 4) /2 so, 2y + 2 = -x + 4 so, x + 2y - 2 = 0.. this the required equation.

OpenStudy (anonymous):

Okay thanks :D

OpenStudy (anonymous):

np..

OpenStudy (anonymous):

@jumboabrfan xD seeeeeeeeeeeeeeeeeeeeeeeeeeee :D

OpenStudy (anonymous):

2x-4y=18 3x-y+10=0 what is the first step

OpenStudy (dan815):



OpenStudy (anonymous):

do any one know

OpenStudy (anonymous):

@ijustdontcare make y the subject of both formulae.. then equate to get x. when u get x, simply put thevalue in any one of the two equations to get y.

OpenStudy (anonymous):

@dan815 ?????????

OpenStudy (anonymous):

don't get it

OpenStudy (anonymous):

2x - 4y = 18 -> y = ( x - 9) / 2 [check on paper] 3x - y + 10 = 0 -> y = 3x + 10 so (x - 9) / 2 = 3x + 10 solve for x. then put value of x in -> y = 3x + 10 to get y.

OpenStudy (dan815):

@xoya wat are you stuck on

OpenStudy (anonymous):

Nothing